OpenStudy (surryyy):

given f(x)=x+3 and g(x)=13x−5 what is f(4)+g(−2) 48 28 -24 24

Nnesha (nnesha):

first find g(-2) and f(4) g(-2) means x=-2 so substitute x for -2 into g(x) function

Nnesha (nnesha):

\[\huge\rm g(\color{reD}{-2})=13\color{reD}{x}-5\]substitute x for -2 then sovle!

Nnesha (nnesha):

solve*

OpenStudy (surryyy):

my final answer would be -24?

Nnesha (nnesha):

are you sure ?

OpenStudy (surryyy):

um yes?

Nnesha (nnesha):

if i ask how would you explain ??

Nnesha (nnesha):

what did you get when u replaced x for -2 into g(x) function ?? :=)

OpenStudy (surryyy):

f(4) + g(-2) = 7 + (-31)

OpenStudy (surryyy):

7 - 31 = -24

Nnesha (nnesha):

thanks :=)